# DiffLight: A Partial Rewards Conditioned Diffusion Model for Traffic Signal Control with Missing Data

### v1

@@ 1 Introduction
To tackle these challenges, we introduce DiffLight, a novel conditional diffusion model for TSC with missing data. We propose a Partial Rewards Conditioned Diffusion (PRCD) model for both traffic data imputation and decision-making under data-missing scenarios to prevent missing rewards from interfering with the learning process. Meanwhile, to effectively capture the spatial-temporal dependencies among intersections, we design the noise model as a Spatial-Temporal transFormer (STFormer) architecture. In addition, we propose a Diffusion Communication Mechanism (DCM) to enable communication and promote the capture of spatiotemporal dependencies in the traffic network through the propagation of generated observations, facilitating better control in scenarios with missing data. Extensive experiments on five datasets with various data-missing scenarios are conducted to evaluate the effectiveness of DiffLight. The experimental results indicate that DiffLight is highly competitive for TSC with missing data.
@@ 2.2 Traffic Signal Control with Missing Data
We formulate TSC with missing data as POMDP, and consider TSC in a traffic network with several intersections. Agents are deployed at each intersection of the traffic network. As illustrated in Figure 1, for a four-way intersection, there are twelve traffic movements from the intersection’s entrance lane $l_{\mathrm{in}}$ to the departure lane $l_{\mathrm{out}}$ and four pairs of traffic movements without conflict comprise four traffic signal phases. For example, the traffic signal phase of the intersection in Figure 1 is phase-A which involves movement-2 =$=\{l_{\mathrm{in}}^{2}\rightarrow l_{\mathrm{out}}^{7},l_{\mathrm{in}}^{2}\rightarrow l_{\mathrm{out}}^{8},l_{\mathrm{in}}^{2}\rightarrow l_{\mathrm{out}}^{9}\}$ and movement-8 =$=\{l_{\mathrm{in}}^{8}\rightarrow l_{\mathrm{out}}^{1},l_{\mathrm{in}}^{8}\rightarrow l_{\mathrm{out}}^{2},l_{\mathrm{in}}^{8}\rightarrow l_{\mathrm{out}}^{3}\}$.
In this paper, the observation $o_{t}$ contains the number of vehicles $L_{\text{num}}$ and queue length $L_{\text{queue}}$ in every entrance lane of the local intersection. Available actions are four phases composed of traffic movements without conflict. The immediate reward $r_{t}$ is defined as the sum of the queue length $\sum_{l_{\mathrm{in}}}{L_{\text{queue}}}$. Due to the lack or error of sensors resulting in missing traffic data, $o_{t}$ and $r_{t}$, which are derived from traffic data, could be missing in a particular pattern. We consider random missing and kriging missing patterns detailed in Appendix C.2. To simplify the problem, we assume that $o_{t}$ and $r_{t}$ in an intersection are missing simultaneously.
@@ 2.3 Diffusion Models for Reinforcement Learning
We make a brief introduction to the diffusion model adopted to address decision-making problems in RL. Among existing works on diffusion-based RL, Diffuser [16] chooses to diffuse on an observation-action trajectory with returns condition to directly generate actions. Decision Diffuser [17] chooses to diffuse only on observation trajectory with returns condition in order to obtain better performance due to the less smooth nature of discrete actions. We focus on introducing the approach diffusing on observation trajectory $\bm{\tau}$ sampled from offline dataset $\mathcal{D}$. We denote the $k$-step denoised output of the diffusion model as $x^{k}(\bm{\tau})$. In this approach, the observation trajectory would be diffused to generate $o_{t+1}$. However, only diffusing on observation trajectory is not enough to make decisions. An inverse dynamics model $f_{\phi}$ is adapted to generate the action $a_{t}$ that makes the observation transit from $o_{t}$ to $o_{t+1}$,
@@
Classifier-free guidance [26] aims to learn the conditional distribution $q({x(\bm{\tau})}|y(\bm{\tau}))$ without a pre-trained classifier. Instead, it learns both a conditional $\epsilon_{\theta}({x}^{k}(\bm{\tau}),y(\bm{\tau}),k)$ and an unconditional $\epsilon_{\theta}({x}^{k}(\bm{\tau}),\phi,k)$ for the noise. Then, the perturbed noise $\epsilon_{\theta}({x}^{k}(\bm{\tau}),\phi,k)+\omega(\epsilon_{\theta}({x}^{k}(\bm{\tau}),y(\bm{\tau}),k)-\epsilon_{\theta}({x}^{k}(\bm{\tau}),\phi,k))$ can be used to generate samples, where $\omega$ is the guidance scale.
@@ 3 Methodology
${y}(\bm{\tau}):=[{r}(\bm{\tau}),{y}^{\prime}(\bm{\tau})],\qquad{y}^{\prime}(\bm{\tau}):=[\bm{\tau}^{\prime}_{\mathrm{obs}},\bm{\tau}^{\prime}_{\mathrm{nei}}],$ (5)
where ${r}(\bm{\tau})$ is observable rewards, $\bm{\tau}^{\prime}_{\mathrm{obs}}$ is the observed part of trajectory $\bm{\tau}$ of local intersection and $\bm{\tau}^{\prime}_{\mathrm{nei}}$ is the observed observations of neighboring intersections. Specifically, the observed part is the historical data collected by sensors, while the observable part refers to the data that could be collected in the past and future. Due to discrete and high-frequent actions in TSC, we choose to diffuse only on observations and use the inverse dynamic model to generate the action, which has a better performance proven in Appendix F.1 and F.6. We define the observation trajectory $\bm{\tau}$ under data-missing scenarios as,
@@
where ${o}_{t}$ is the observation that has been observed, $\hat{o}_{t}$ is the unobserved observation that needs to be generated, $C$ is the length of historical observations and $H$ is the horizon of future observations. According to the definition of observation trajectory $\bm{\tau}$, we specify $\bm{\tau}^{\prime}_{\mathrm{obs}}$ as the observed part of $\bm{\tau}$ from time $t-C+1$ to $t$. Additionally, $\bm{\tau}^{\prime}_{\mathrm{nei}}=\cup_{N}f_{\mathrm{nei}}(\bm{\tau}^{i})$ represents the observations of entrance lanes from all neighboring intersections that feed into the entrance lanes of the local intersection, as shown in Figure 1, where $\bm{\tau}^{i}$ denotes the observation trajectory of the neighboring intersection $i$ and $N$ is total number of neighboring intersections. Furthermore, we define $\bm{\tau}_{\mathrm{obs}}$ as the observable part of the entire trajectory and $\bm{\tau}_{\mathrm{mis}}$ as the missing part.We detail ${r}(\bm{\tau})$ in Section 3.1.
@@
Figure 2: An overview of DiffLight. We demonstrate the signal control process of an intersection in random missing. Traffic data is collected by sensors to derive rewards and observations. Missing rewards and observations are masked. Only the observed part of the observation trajectory and observable rewards from the local intersection and observation trajectories of neighboring intersections would be input into PRCD with STFormer. In the inference process, DCM would work with STFormer to generate observations. The inverse dynamics model is used to generate the action to control the traffic signal.
@@ 3.1 Partial Rewards Conditioned Diffusion
We adopt partial rewards as the condition instead of returns [16, 17, 19], which can be expressed as,
@@
where ${r}_{t}$ is the observable reward and $\tilde{r}_{t}$ is the missing reward. Since under data-missing scenarios, sensors deployed to collect traffic data for rewards may malfunction or be lacking. The absence of partial rewards makes it hard to calculate returns. Therefore, we have to explore how to make decisions with partial observable rewards.
@@
$p_{\theta}({x}^{0}(\bm{\tau})|{y}(\bm{\tau}))=p_{\theta}({x}^{0}(\bm{\tau}_{\mathrm{obs}})|{r}(\bm{\tau}),{y}^{\prime}(\bm{\tau}))\cdot p_{\theta}({x}^{0}(\bm{\tau}_{\mathrm{mis}})|{y}^{\prime}(\bm{\tau}))\\ .$ (8)
We parameterize Equation 3 as the same form of locally conditioned diffusion [27, 28], propose partial rewards conditioned diffusion (PRCD) with classifier-free guidance [26] and introduce it into TSC with missing data. Given condition set $\mathbf{c}=\{\phi,{r}(\bm{\tau})\}$ and binary non-overlapping mask set $\mathbf{m}=\{m_{\mathrm{mis}},m_{\mathrm{obs}}\}$, PRCD assigns partial rewards to corresponding observation sub-trajectory masked by $\mathbf{m}$, which can be formulated as,
$\begin{split}\epsilon_{\theta}({x}^{k}(\bm{\tau}),k,{y}^{\prime}(\bm{\tau}),\mathbf{c},\mathbf{m}):=&m_{\mathrm{mis}}\odot{\epsilon}_{\theta}({x}^{k}(\bm{\tau}),k,{y}^{\prime}(\bm{\tau}),\phi)+\\ &m_{\mathrm{obs}}\odot{\epsilon}_{\theta}({x}^{k}(\bm{\tau}),k,{y}^{\prime}(\bm{\tau}),{r}(\bm{\tau})),\\ {\epsilon}_{\theta}({x}^{k}(\bm{\tau}),k,{y}^{\prime}(\bm{\tau}),c_{i}):=&\epsilon_{\theta}({x}^{k}(\bm{\tau}),k,{y}^{\prime}(\bm{\tau}),\phi)+\\ &\alpha(\epsilon_{\theta}({x}^{k}(\bm{\tau}),k,{y}^{\prime}(\bm{\tau}),c_{i})-\epsilon_{\theta}({x}^{k}(\bm{\tau}),k,{y}^{\prime}(\bm{\tau}),\phi)),\end{split}$ (9)
@@ 3.3 Diffusion Communication Mechanism
$\bm{\tau}^{\prime}_{\mathrm{nei}}=\left\{\begin{aligned} &\cup_{N}f_{\mathrm{nei}}(\bm{\tau}^{i}),&k=K,\\ &\cup_{N}f_{\mathrm{nei}}(\frac{1}{\sqrt{\bar{\alpha}^{k}}}({x}^{k}(\bm{\tau}^{i})-\sqrt{1-\bar{\alpha}^{k}}{\epsilon}_{\theta}({x}^{k}(\bm{\tau}),k,{y}^{\prime}(\bm{\tau}),\mathbf{c},\mathbf{m}))),&k<K.\end{aligned}\ \right.$ (10)
@@ 3.4 Training and Inference of DiffLight
$\begin{split}\mathcal{L}(\theta):=&\mathbb{E}_{(o,a,o^{\prime})\in\mathcal{D}}[\|a-f_{\phi}(o,o^{\prime})\|^{2}\cdot\mathds{1}(o,o^{\prime})]+\\ &\mathbb{E}_{k,\epsilon,\beta\sim\mathrm{Bern}(p)}[\|\epsilon_{\theta}({x}^{k}(\bm{\tau}),k,{y}^{\prime}(\bm{\tau}),(1-\beta){r}(\bm{\tau})+\beta\phi)-\epsilon\|^{2}].\end{split}$ (11)
For each trajectory $\bm{\tau}$, we sample noise $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and a diffusion timestep $k\sim\mathcal{U}(1,K)$, construct a masked noise array of observations ${x}^{k}(\bm{\tau})$ with $\bm{\tau}^{\prime}_{\mathrm{obs}}$, and predict the noise $\hat{\epsilon}_{\theta}({x}^{k}(\bm{\tau}),k,{y}^{\prime}(\bm{\tau}),{r}(\bm{\tau}))$. Missing values in condition $\bm{\tau}^{\prime}_{\mathrm{nei}}$ are padded with zeros. It should be noted that we ignore the rewards condition ${r}(\bm{\tau})$ with probability $p$ and the inverse dynamics is trained with individual transitions without missing observation $o$ or $o^{\prime}$. For the training process of DiffLight, due to the inaccessibility of the ground-truth of missing data, we consider it self-supervised learning. In the pattern of random missing, given a trajectory $\bm{\tau}$ and conditions, we can separate the observed part into two parts and set one of them to miss. In the pattern of kriging missing, we randomly mask the whole trajectories of one observed intersection. Then, we can train the noise model $\epsilon_{\theta}$ by solving Equation 11.
@@
DiffLight is deployed to every intersection in the traffic network in the inference process. Given rewards ${r}(\bm{\tau})$, a $C$-length observed trajectory of local intersection $\bm{\tau}^{\prime}_{\mathrm{obs}}$ with missing data and trajectories of neighboring intersections $\bm{\tau}^{\prime}_{\mathrm{nei}}$, the agent can impute the missing observations of local intersection, predict the observations in the future and generate next action with Equation 2 and 9. In order to sample a high reward trajectory, the rewards from $t+1$ to $t+H$ are considered in ${r}(\bm{\tau})$, which is set to 1.

### v2

@@ 1 Introduction
To tackle these challenges, we introduce DiffLight, a novel conditional diffusion model for TSC with missing data. We propose a Partial Rewards Conditioned Diffusion (PRCD) model for both traffic data imputation and decision-making under data-missing scenarios to prevent missing rewards from interfering with the learning process. Meanwhile, to effectively capture the spatial-temporal dependencies among intersections, we design the noise model as a Spatial-Temporal transFormer (STFormer) architecture. In addition, we propose a Diffusion Communication Mechanism (DCM) to enable communication and promote the capture of spatio-temporal dependencies in the traffic network through the propagation of generated observations, facilitating better control in scenarios with missing data. Extensive experiments on five datasets with various data-missing scenarios are conducted to evaluate the effectiveness of DiffLight. The experimental results indicate that DiffLight is highly competitive for TSC with missing data.
@@ 2.2 Traffic Signal Control with Missing Data
We formulate TSC with missing data as POMDP, and consider TSC in a traffic network with several intersections. Agents are deployed at each intersection of the traffic network. As illustrated in Figure 1, for a four-way intersection, there are twelve traffic movements from the intersection’s entrance lane $l_{\mathrm{in}}$ to the departure lane $l_{\mathrm{out}}$, and four pairs of traffic movements without conflict comprise four traffic signal phases, i.e., A, B, C, D. For example, the traffic signal phase of the intersection in Figure 1 is phase-A which involves movement-2 =$=\{l_{\mathrm{in}}^{2}\rightarrow l_{\mathrm{out}}^{7},l_{\mathrm{in}}^{2}\rightarrow l_{\mathrm{out}}^{8},l_{\mathrm{in}}^{2}\rightarrow l_{\mathrm{out}}^{9}\}$ and movement-8 =$=\{l_{\mathrm{in}}^{8}\rightarrow l_{\mathrm{out}}^{1},l_{\mathrm{in}}^{8}\rightarrow l_{\mathrm{out}}^{2},l_{\mathrm{in}}^{8}\rightarrow l_{\mathrm{out}}^{3}\}$.
In this paper, the observation $o_{t}$ contains the number of vehicles $L_{\text{num}}$ and queue length $L_{\text{queue}}$ in every entrance lane of the local intersection. Available actions are four phases. The immediate reward $r_{t}$ is defined as the sum of the queue length $\sum_{l_{\mathrm{in}}}{L_{\text{queue}}}$. Due to the lack or error of sensors resulting in missing traffic data, $o_{t}$ and $r_{t}$, which are derived from traffic data, could be missing in a particular missing pattern. We consider random missing and kriging missing patterns detailed in Appendix C.2. To simplify the problem, we assume that $o_{t}$ and $r_{t}$ in an intersection are missing simultaneously.
@@ 2.3 Diffusion Models for Reinforcement Learning
Recently, many diffusion-based approaches have been proposed to address decision-making problems in RL. Among existing works, Diffuser [16] chooses to diffuse on an observation-action trajectory with returns as the condition to generate actions directly. Decision Diffuser [16] focuses solely on diffusing the observation trajectory conditioned on returns. By avoiding direct diffusion over actions, Decision Diffuser enhances performance in scenarios with discrete actions. In this paper, considering the discrete nature of actions in TSC, we focus on introducing the approach diffusing on the observation trajectory $\bm{\tau}$ sampled from offline dataset $\mathcal{D}$. We denote the $k$-step denoised output of the diffusion model as $x^{k}(\bm{\tau})$. The observation trajectory would be diffused to generate $o_{t+1}$. However, only diffusing on observation trajectory is not enough to make decisions. An inverse dynamics model $f_{\phi}$ is adapted to generate the action $a_{t}$ that makes the observation transit from $o_{t}$ to $o_{t+1}$,
@@
Classifier-free guidance [26] aims to learn the conditional distribution $q({x(\bm{\tau})}|y(\bm{\tau}))$. It learns both a conditional $\epsilon_{\theta}({x}^{k}(\bm{\tau}),y(\bm{\tau}),k)$ and an unconditional $\epsilon_{\theta}({x}^{k}(\bm{\tau}),\phi,k)$ for the noise. Then, the perturbed noise $\epsilon_{\theta}({x}^{k}(\bm{\tau}),\phi,k)+\omega(\epsilon_{\theta}({x}^{k}(\bm{\tau}),y(\bm{\tau}),k)-\epsilon_{\theta}({x}^{k}(\bm{\tau}),\phi,k))$ can be used to generate samples, where $\omega$ is the guidance scale.
@@ 3 Methodology
${y}(\bm{\tau}):=[{r}(\bm{\tau}),{y}^{\prime}(\bm{\tau})],\qquad{y}^{\prime}(\bm{\tau}):=[\bm{\tau}^{\prime}_{\mathrm{obs}},\bm{\tau}^{\prime}_{\mathrm{nei}}].$ (5)
To facilitate a better understanding of the symbol definitions, we categorize the trajectory segments into three types. The observed trajectory consists of data collected by sensors up to time $t$. The missing trajectory includes data that have not been collected during this period. The observable trajectory encompasses both the observed trajectory and the data that could potentially be collected in the future. In this context, ${r}(\bm{\tau})$ is the observable reward trajectory, and $\bm{\tau}^{\prime}_{\mathrm{obs}}$ is the observed part of trajectory $\bm{\tau}$ from the local intersection. $\bm{\tau}^{\prime}_{\mathrm{nei}}=\cup_{N}f_{\mathrm{nei}}(\bm{\tau}^{i})$ is the observed observations from neighboring intersections, where $\bm{\tau}^{i}$ denotes the observation trajectory of the neighboring intersection $i$, $N$ is total number of neighboring intersections, $f_{\mathrm{nei}}(\cdot)$ represents the observed observations from entrance lanes of all neighboring intersections that feed into the entrance lanes of the local intersection, as shown in Figure 1. Due to discrete and high-frequent actions in TSC, we choose to diffuse solely on observations and utilize the inverse dynamic model to generate actions, which demonstrates a better performance proven in Appendix F.1 and F.6. We define the observation trajectory $\bm{\tau}$ under data-missing scenarios as,
@@
where ${o}_{t}$ is the observation collected by sensors, $\hat{o}_{t}$ is the uncollected observation up to time $t$ or a potentially collectible observation in the future, $C$ is the length of historical observations and $H$ is the horizon of future observations. It should be noted that we generate $H$-step future observations to enable effective long-horizon planning [16].
@@
Figure 2: An overview of DiffLight. We demonstrate the signal control process of an intersection in random missing. Traffic data is collected by sensors to derive rewards and observations. Missing rewards and observations are masked. Only the observed part of the observation trajectory and observable rewards from the local intersection, and observation trajectories from neighboring intersections would be input into PRCD with STFormer. In the inference process, DCM would work with STFormer to generate observations. The inverse dynamics model is used to generate actions to control the traffic signal.
@@ 3.1 Partial Rewards Conditioned Diffusion
Under data-missing scenarios, sensors deployed to collect traffic data for rewards may malfunction or be lacking. The absence of partial rewards makes it challenging to calculate returns. Therefore, we adopt partial rewards as the condition instead of returns [16, 17, 19], which can be expressed as,
@@
where ${r}_{t}$ is the collected reward or a potentially collectible reward in the future, and $\tilde{r}_{t}$ is the uncollected reward.
@@
$p_{\theta}({x}^{0}(\bm{\tau})|{y}(\bm{\tau}))=p_{\theta}({x}^{0}(\bm{\tau}_{\mathrm{obs}})|{r}(\bm{\tau}),{y}^{\prime}(\bm{\tau}))\cdot p_{\theta}({x}^{0}(\bm{\tau}_{\mathrm{mis}})|{y}^{\prime}(\bm{\tau})),$ (8)
where $\bm{\tau}_{\mathrm{obs}}$ is the observable part of the trajectory $\bm{\tau}$ from the local intersection, and $\bm{\tau}_{\mathrm{mis}}$ is the missing part. We parameterize Equation 8 as the same form of locally conditioned diffusion [27, 28], propose partial rewards conditioned diffusion (PRCD) with classifier-free guidance [26] and introduce it into TSC with missing data. Given condition set $\mathbf{c}=\{\phi,{r}(\bm{\tau})\}$ and binary non-overlapping mask set $\mathbf{m}=\{m_{\mathrm{mis}},m_{\mathrm{obs}}\}$, PRCD assigns partial rewards to corresponding observation sub-trajectory masked by $\mathbf{m}$, which can be formulated as,
$\begin{split}\hat{\epsilon}_{\theta}({x}^{k}(\bm{\tau}),k,{y}^{\prime}(\bm{\tau}),\mathbf{c},\mathbf{m}):=&m_{\mathrm{obs}}\odot\hat{\epsilon}_{\theta}({x}^{k}(\bm{\tau}),k,{y}^{\prime}(\bm{\tau}),{r}(\bm{\tau}))+\\ &m_{\mathrm{mis}}\odot\hat{\epsilon}_{\theta}({x}^{k}(\bm{\tau}),k,{y}^{\prime}(\bm{\tau}),\phi),\end{split}$ (9)
where $\hat{\epsilon}_{\theta}({x}^{k}(\bm{\tau}),k,{y}^{\prime}(\bm{\tau}),c_{i})$ could be expressed using classifier-free guidance,
$\begin{split}\hat{\epsilon}_{\theta}({x}^{k}(\bm{\tau}),k,{y}^{\prime}(\bm{\tau}),c_{i}):=&\epsilon_{\theta}({x}^{k}(\bm{\tau}),k,{y}^{\prime}(\bm{\tau}),\phi)+\\ &\omega(\epsilon_{\theta}({x}^{k}(\bm{\tau}),k,{y}^{\prime}(\bm{\tau}),c_{i})-\epsilon_{\theta}({x}^{k}(\bm{\tau}),k,{y}^{\prime}(\bm{\tau}),\phi)),\end{split}$ (10)
@@ 3.3 Diffusion Communication Mechanism
$\bm{\tau}^{\prime}_{\mathrm{nei}}=\left\{\begin{aligned} &\cup_{N}f_{\mathrm{nei}}(\bm{\tau}^{i}),&k=K,\\ &\cup_{N}f_{\mathrm{nei}}(\frac{1}{\sqrt{\bar{\alpha}^{k}}}({x}^{k}(\bm{\tau}^{i})-\sqrt{1-\bar{\alpha}^{k}}\hat{\epsilon}_{\theta}({x}^{k}(\bm{\tau}),k,{y}^{\prime}(\bm{\tau}),\mathbf{c},\mathbf{m}))),&k<K.\end{aligned}\ \right.$ (11)
@@ 3.4 Training and Inference of DiffLight
$\begin{split}\mathcal{L}(\theta):=&\mathbb{E}_{(o,a,o^{\prime})\in\mathcal{D}}[\|a-f_{\phi}(o,o^{\prime})\|^{2}\cdot\mathds{1}(o,o^{\prime})]+\\ &\mathbb{E}_{k,\epsilon,\beta\sim\mathrm{Bern}(p)}[\|\hat{\epsilon}_{\theta}({x}^{k}(\bm{\tau}),k,{y}^{\prime}(\bm{\tau}),(1-\beta){r}(\bm{\tau})+\beta\phi,\mathbf{m})-\epsilon\|^{2}].\end{split}$ (12)
For each trajectory $\bm{\tau}$, we sample noise $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and a diffusion timestep $k\sim\mathcal{U}(1,K)$, construct a masked noise array of observations ${x}^{k}(\bm{\tau})$ with $\bm{\tau}^{\prime}_{\mathrm{obs}}$, and predict the noise $\hat{\epsilon}_{\theta}({x}^{k}(\bm{\tau}),k,{y}^{\prime}(\bm{\tau}),{r}(\bm{\tau}),\mathbf{m})$. Missing values in condition $\bm{\tau}^{\prime}_{\mathrm{nei}}$ are padded with zeros. It should be noted that we ignore the rewards condition ${r}(\bm{\tau})$ with probability $p$ and the inverse dynamics is trained with individual transitions without missing observation $o$ or $o^{\prime}$. For the training process of DiffLight, due to the inaccessibility of the ground-truth of missing data, we consider it self-supervised learning. In random missing, given a trajectory $\bm{\tau}$ and conditions, we can separate the observed part into two parts and set one of them to miss. In kriging missing, we randomly mask the whole trajectories of one observed intersection. Then, we can train the noise model $\hat{\epsilon}_{\theta}$ by solving Equation 12.
@@
DiffLight is deployed to every intersection in the traffic network in the inference process. Given rewards ${r}(\bm{\tau})$, a $C$-step observed trajectory of local intersection $\bm{\tau}^{\prime}_{\mathrm{obs}}$ with missing data and trajectories of neighboring intersections $\bm{\tau}^{\prime}_{\mathrm{nei}}$, the agent can impute the missing observations of local intersection, predict the observations in the future and generate next action with Equation 2, 9 and 10. In order to sample a high reward trajectory, the rewards from time $t+1$ to $t+H$ are considered in ${r}(\bm{\tau})$, which is set to 1.